### v1

# From Freshness to Effectiveness: Goal-Oriented Sampling for Remote Decision Making

Abstract
Data freshness, measured by Age of Information (AoI), is highly relevant in networked applications such as Vehicle to Everything (V2X), smart health systems, and Industrial Internet of Things (IIoT). Yet, freshness alone does not equate to informativeness. In decision-critical settings, some stale data may prove more valuable than fresh updates. To explore this nuance, we move beyond AoI-centric policies and investigate how data staleness impacts decision-making under data-staleness-induced uncertainty. We pose a central question: What is the value of information, when freshness fades—and only its power to shape remote decisions remains? To capture this endured value, we propose AR-MDP, an Age-aware Remote Markov Decision Process framework, which co-designs optimal sampling and remote decision-making under a sampling frequency constraint and random delay. To efficiently solve this problem, we design a new two-stage hierarchical algorithm namely Quick Bellman-Linear-Program (QuickBLP), where the first stage involves solving the Dinkelbach root of a Bellman variant and the second stage involves solving a streamlined linear program (LP). For the tricky first stage, we propose a new One-layer Primal-Dinkelbach Synchronous Iteration (OnePDSI) method, which overcomes the re-convergence and non-expansive divergence present in existing per-sample multi-layer algorithms. Through rigorous convergence analysis of our proposed algorithms, we establish that the worst-case optimality gap in OnePDSI exhibits exponential decay with respect to iteration $K$ at a rate of $\mathcal{O}(\frac{1}{R^{K}})$. Through sensitivity analysis, we derive a threshold for the sampling frequency, beyond which additional sampling does not yield further gains in decision-making. Simulation results validate our analyses.
Index Terms:
Age of Information, Value of Information, Markov Decision Process, Remote Decision Making, Goal-Oriented Communications, Effective Communications
I Introduction
Age of Information (AoI) is a crucial metric for evaluating information freshness in status update systems, garnering broad attention from both academia and industry [2, 3]. Currently, AoI has been applied in a wide range of applications such as queue control [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], source coding[14, 15], remote estimation [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], and network design [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] (see [39] for a comprehensive review). Central to this field is the question: “How can we minimize the Age of Information?” The conventional wisdom in AoI optimization lies in an intuitively compelling yet mathematically non-trivial heuristic: “fresher information holds greater value”. This heuristic finds validation across real-world applications. In Internet of Vehicles (IoV) systems, timely status updates are essential for enabling safety-critical driving maneuvers. In financial markets, access to first-hand information directly impacts the effectiveness of trading decisions. These examples empirically demonstrate that minimizing AoI can improve estimation accuracy or enhance subsequent information-driven decision-making.
A significant challenge in the field is the scarcity of rigorous analytical frameworks that establish clear relationships between information freshness and value. In many scenarios, freshness alone does not determine how valuable a data update is to the eventual decision it informs. Instead, the value of information may depend on multiple interrelated factors beyond AoI, including the semantic content of transmitted packets and the underlying dynamics of the monitored source[40, 41, 42, 43]. This recognition has driven the development of various AoI variants. One approach introduces nonlinear AoI penalties, implemented through both empirical configurations [44, 45, 46, 47] and theoretically derived functions [48, 49, 50, 51]. These nonlinear formulations aim to quantify the loss resulting from information staleness. Additionally, researchers have proposed various heuristic metrics, including Age of Synchronization (AoS) [52], Age of Incorrect Information (AoII) [53], and Age of Changed Information (AoCI) [54], and Age of Collected Information [55, 56]. These metrics customize time-related penalization from a wider perspective than what can be captured with age, particularly in applications involving rapidly evolving source dynamics. For remote estimation, mean square estimation error (MSEE) [16] and context-aware Urgency of Information (UoI) [57] are leveraged to penalize the real-time reconstruction distortion. Although information freshness is widely regarded as beneficial in communication systems, its specific value to decision quality has not been fully characterized. This leads to a key question: From the perspective of remote decision-making, what kind of information truly matters, and how should its value guide what is sampled and transmitted? This open question motivates our theoretical investigation in this paper.
Several works have proposed heuristic approaches to characterize this relationship. In [58], Dong et al. introduced Age upon Decisions (AuD), which measures the time elapsed between data generation and its use in decision-making, where the decision epoch follows a stochastic distribution. In a similar vein, [59] proposed Age of Actuation (AoA). In [60], three types of decisions: correct decisions, incorrect decisions, and missed decisions are assigned different time-cost functions. A new metric termed Penalty upon Decision (PuD) was proposed. In [41] and [61], a tensor-based metric termed Goal-oriented Tensor (GoT) was proposed as a unified framework for existing metrics. However, while prior work has advanced communication optimization through a variety of metrics, it often overlooks how decision-making systems actually function when operating with potentially stale information. The core challenge is to bridge the gap between metric-driven optimization and the realities of decision-making under data-staleness-induced uncertainty.
In this paper, we aim to examine how stale information impacts remote stochastic decision-making under conditions of data-staleness-induced uncertainty. To this end, we propose the Age-aware Remote Markov Decision Process [^1] In [62], the term remote MDP was first proposed as a pathway to pragmatic or goal-oriented communications. Our paper focuses on the communication delay and introduces the age to enhance remote decision-making to achieve a certain goal, hence the term age-aware remote MDP. (AR-MDP), a comprehensive framework that jointly optimizes sampling and sequential decision-making, with a specific purpose to achieve goal-oriented effective decisions. The relationship between sampling and decision-making exhibits inherent bidirectional coupling. Sampling policies directly influence the staleness-induced uncertainty, which can lead to unsatisfactory decision outcomes. Conversely, decision-making processes affect the stochastic evolution of the source system, which in turn impacts the effectiveness of content-driven goal-oriented sampling mechanisms. To decouple these two processes and achieve optimal decision making under random delay and sampling frequency constraint, we formulate the problem as a constrained Partially Observable Semi-Markov Decision Process problem, where AoI no longer serves as a typical indicator, but instead plays a critical role as side information for enhancing remote, goal-oriented decisions.. We design efficient algorithms to solve this problem. To the best of our knowledge, this is the first work that introduces AoI into the time-lag MDP family, and the first work that explores AoI’s new role as side information to facilitate remote decision making under random delay.
II Related Work and Our Novelty
II-A Sampling Under Random Delay
The results in this paper contribute to the optimal sampling design under random delay. In real-world network environments, communication channels inevitably experience random delays due to various factors: network handover, congestion, variable sample sizes, and packet retransmissions. These fundamental characteristics have driven research into developing optimal sampling policies under random delay. Existing literature has focused on optimizing three key aspects: $i$) information freshness [63, 64, 65, 66, 67, 68]; $ii$) remote estimation [25, 16, 17, 69, 70, 71]; and $iii$) remote inference [72, 73, 74, 75] under random delay. A particularly noteworthy and counter-intuitive finding in this field reveals that optimal sampling may require the source to deliberately wait before submitting a new sample to the channel, challenging the conventional wisdom of throughput-optimal zero-wait sampling policy.
Information Freshness: In the seminal work [63], Sun et al. derived an AoI-optimal sampling policy under random delay. This paper revealed that under a maximum rate constraint, the AoI-optimal sampling follows a threshold structure, where sampling is activated only when the current AoI exceeds a specific threshold determinable through a low-complexity bisection search method. In [64], the optimal sampling policy for a non-linear monotonic function of AoI was designed. Tang et al. [65] extended this framework to scenarios with unknown delay statistics, employing stochastic approximation methods to determine the AoI-optimal sampling policy under unknown delay statistics. Further advancements were made by Pan et al. [66], who developed AoI-optimal sampling policies under unreliable transmission with random two-way delay. Most recently, Liyanaarachchi and Ulukus enhanced [63] by incorporating random ACK delay, demonstrating that sampling before receiving acknowledgment can potentially achieve superior AoI performance [67]. In [68], Peng et al. designed optimal sampling policies that achieve minimal Age of Changed Information (AoCI) [76]—a metric capable of detecting source changes—under known and unknown delay statistics respectively.
Remote Estimation: The theoretical foundations of remote estimation-oriented sampling under random delay were established through [16, 17]. These studies developed an optimal sampling policy for the Wiener process that minimizes the mean square estimation error (MSEE) while adhering to sampling frequency constraints. Their research revealed that the optimal sampling policy exhibits a threshold structure, where sampling is initiated only when the real-time MSEE surpasses a predetermined threshold. Building upon this foundation, Ornee et al. [25] expanded this theoretical framework by investigating MSEE-optimal sampling for the Ornstein-Uhlenbeck (OU) process, a stationary Gauss-Markov process, under random delay. In [69] and [70], the MSEE-optimal sampling policies are derived for the Wiener process and the OU process under unknown delay statistics. In [71], Chen et al. derived the optimal sampling policy that achieves minimized uncertainty-of-information under random delay, where UoI is defined as the conditional entropy of the source at the receiver given the observation history [51]—mathematically expressed as $H(X_{t}|\mathcal{H}_{t})$, with $X(t)$ representing the source state at time $t$ and $\mathcal{H}_{t}$ denoting the available observation history at the receiver. Most recently, Chen et al. [77] derived AoII-optimal sampling policies under random delay.
Remote Inference: Recent research has revealed insights into remote inference performance and its relationship with information freshness metrics. In [73, 72], Shisher et al. demonstrated that the loss function in remote inference may not be monotonic in terms of the age of the samples (features) used, if the source sequence is not Markovian. Upon making this remarkable observation, the authors developed policies that allow selection of aged samples from the buffer, rather than the freshest one. This was termed the “selection-from-buffer” model. In [74], a learning and communication co-design problem that jointly optimizes feature length selection and transmission scheduling is proposed. In [75], Ari et al expanded previous works by incorporating time-varying statistics of random delay and delayed feedback, developing optimal sampling policies to minimize long-term inference error within the “selection-from-buffer” model.
To the best of our knowledge, the optimal sampling policy for Remote Decision Making under random delay remains an open research problem, which we address in this paper.
TABLE I: Comparisons of Time-Lag MDPs
Type Observation Reference
Standard MDP $O(t)=X_{t}$ [78]
DDMDP $O(t)=X_{t-d}$ [79]
SDMDP $O(t)=X_{t-D}$ [80]
Age-Aware Remote MDP $O(t)=X_{t-\Delta(t)}$ This Work
Figure 1: Comparisons among standard MDP, DDMDP, SDMDP, and AR-MDP.
II-B Decision-Making over Stale Status
The proposed AR-MDP in this paper also enriches the family of time-lag MDP, whose focus is on making decisions based on stale status. As illustrated in Table I and Fig. 1, two primary types of MDPs address observation delay at the decision maker: deterministic delayed MDP (DDMDP) [79] and stochastic delayed MDP (SDMDP) [80]. The DDMDP introduces a constant observation delay $d$ to the standard MDP framework. At any given time $t$, the decision-maker accesses the time-varying data as $O(t)=X_{t-d}$. The main result of the DDMDP problem is its reducibility to a standard MDP without delays through state augmentation, as detailed by Altman and Nain [79]. The SDMDP extends DDMDP by treating the observation delay not as a static constant but as a random variable $D$ following a given distribution $\Pr(D=d)$, with $O(t)=X_{t-D}$. In 2003, V. Katsikopoulos and E. Engelbrecht showed that an SDMDP is also reducible to a standard MDP problem without delay [80]. Thus, it becomes clear to solve an SDMDP problem by solving its equivalent standard MDP.
Figure 2: Evolution of the age $\Delta(t)$ over time.
However, the above time-lag MDPs, where the observation delay follows a given distribution (DDMDP can be regarded as a special type of SDMDP), potentially assume that the state is sampled and transmitted to the decision maker at every time slot [^2] In this case, each state $X_{i},\forall i\in\{0,1,...n\}$ are all sampled and forwarded to the decision maker. The observation delay $D$ is i.i.d random variable and is independent of the sampling policy.. This setup presumes that the system can transmit every state update without encountering any backlog. In practice, constantly sampling and transmitting may result in infinitely accumulated packets in the queue, resulting in severe congestion. This motivates the need for queue control and adaptive sampling policy design in the network [81, 63, 22, 65, 82, 83], where Age of Information (AoI) serves as a key performance indicator. Suppose the $i$-th sample is generated at time $S_{i}$ and is delivered at the receiver at time $D_{i}$. Then, AoI is defined as a sawtooth piecewise function:
$\Delta(t)=t-S_{i},D_{i}\leq t<D_{i+1},\forall i\in\mathbb{N},$ (1)
as shown in Fig. 2. From this definition, the most recently available information at the receiver at time slot $t$ is $O(t)=X_{t-\Delta(t)}$. Different from the DDMDP and SDMDP where the time lag is a constant $d$ or an i.i.d random variable $D$, with $O(t)=X_{t-d}$ or $O(t)=X_{t-D}$, the time lag in the practical network with queue and finite transmission rate is a controlled stochastic process $\Delta(t)$, with $O(t)=X_{t-\Delta(t)}$.
To the best of our knowledge, the design of optimal remote decision making under controlled stochastic delay remains an unexplored research direction, which we address in this paper by studying AR-MDP.
II-C The Novelty of Our Work
  • System Model: This paper proposes AR-MDP, a novel theoretical framework integrating optimal sampling and decision-making under random delay. Differing from prior sampling designs which often treat information as an end in itself—optimizing for freshness, accuracy, or estimation quality under random delays [63, 64, 65, 66, 67, 68, 72, 73, 74, 75, 25, 16, 17, 69, 70, 71], we treat information as a means to action, where its value is defined not by how precise it is, but by how well it enables timely and effective decisions under uncertainty and delay. By modeling the controlled stochastic “delay” as the AoI $\Delta(t)$, AR-MDP also generalizes classical time-lag MDP theoretical framework [79, 80] into a practically relevant model. Through sensitivity analysis, we uncover a fundamental sampling frequency threshold. Beyond this threshold, further increasing the sampling frequency yields no improvement in decision performance. This finding reveals an information-theoretic limit: in decision-centric systems under random delay, the value of information is not solely determined by its freshness, but by its timely relevance to control.
  • Methodology Design: We design QuickBLP, a computationally efficient single-layer algorithm that addresses limitations associated with iterative re-convergence common in multi-layer per-sample algorithms, e.g., [71, Algorithm 1], [84, Algorithm 1], and [85, Section IV.C]. This algorithm is designed based on a key analytical insight that the optimal solution exhibits a threshold structure and can be obtained through a two-stage process. The first stage determines the Dinkelbach root of a Bellman variant, for which we develop OnePDSI, a Cauchy sequence that converges asymptotically to the root without requiring re-convergence. The second stage involves finding the Dinkelbach root of a per-sample constrained Markov Decision Process (CMDP), which traditionally necessitates multiple CMDP solutions and suffers from re-convergence. We resolve this challenge by proving that the Dinkelbach root can be explicitly calculated through the optimal value of a linear program (LP), enabling direct solution to the root through a single LP solution. To the best of our knowledge, QuickBLP is the first framework to tackle constrained partially observable SMDPs through a streamlined single-layer flow, fundamentally improving efficiency by eliminating re-convergence overhead.
  • Theoretical Rigor and Convergence: We significantly advance our previous work [1] by resolving critical convergence limitations. To overcome inherent divergence risks, we design two novel algorithms in this paper: Bisec-$\tau$-RVI and OnePDSI. Although convergence analyses often rely on the Banach Contraction Mapping Theorem [86, Theorem 6] [66, 85], this method falls short in capturing the behavior of our models. Our approach departs from this tradition, providing rigorous proofs that both algorithms guarantee efficient exponential convergence in worst-case settings. These guarantees reinforce the reliability and efficiency of our algorithms in practical applications.
III System Model and Problem Formulation
We consider a time-slotted age-aware remote MDP problem illustrated in Fig. 1(c). Let $X_{t}\in\mathcal{S}$ be the controlled source of interest at time slot $t$. The evolution of the source is a Markov decision process, characterized by the transition probability $\Pr(X_{t+1}|X_{t},a_{t})$ [^3] For short-hand notations, we use the transition probability matrix $\mathbf{P}_{a}$ to encapsulate the dynamics of the source given an action $a_{t}=a$., where $a_{t}\in\mathcal{A}$ represents the controlled action taken by the remote decision maker to control the source in the desired way. The sampler conducts the sampling action $u_{t}\in\{0,1\}$, with $u_{t}=1$ representing the sampling action and $u_{t}=0$ otherwise. Let $S_{i}$ be the sampling time of the $i$-th delivered packet, and $D_{i}$ be the corresponding delivery slot. The random channel delay of the $i$-th packet is denoted as $Y_{i}\in\mathcal{Y}\subseteq\mathbb{N}^{+}$, which is independent of the source $X_{t}$ and is bounded $\max[{Y_{i}}]<\infty$. The sampling times $S_{0},S_{1},\cdots$ shown in Fig. 2 record the time stamp when $u_{t}=1$, given by
$S_{i}=\max\{t\left|t\leq D_{i},u_{t}=1\right.\},\forall i\in\mathbb{N},$ (2)
where the initial state of the system is $S_{0}=0$ and $\Delta(0)=\Delta_{0}$.
At the sampling time $S_{i},\forall i\in\mathbb{N}$, the state $X_{S_{i}}$ along with the corresponding time stamp $S_{i}$ is encapsulated into a packet $(X_{S_{i}},S_{i})$, which is transmitted to a remote decision maker. Upon the receipt of the packet $(X_{S_{i}},S_{i})$ during $t\in[D_{i},D_{i+1})$, the observation history at the decision maker is $\{(X_{S_{j}},S_{j}):{j\leq i}\}$. By employing (1), this sequence can be equivalently represented as $\{(X_{k-\Delta(k)},\Delta(k)):{k\leq t}\}$.
Let $\mathcal{I}_{t}=\{(X_{k-\Delta(k)},\Delta(k),a_{k-1}^{S},a_{k-1}):{k\leq t}\}$ denote the observation history and action history (or simply history) available to the decision maker up to time $t$. The decision maker is tasked with determining both the sampling action and the controlled action $(u_{t},a_{t})\in\{0,1\}\times\mathcal{A}$ at each time slot $t$ by leveraging the history $\mathcal{I}_{t}$. A decision policy of the decision maker is defined as a mapping from the history to a distribution over the joint action space $\{0,1\}\times\mathcal{A}$, denoted by
$\pi_{t}:\mathcal{I}_{t}\rightarrow\mathcal{P}(\mathcal{A}\times\{0,1\}),$ (3)
where $\mathcal{P}(\cdot)$ is a simplex space which represents the probability that an action is taken. Similar to [63], we assume that the sampling policy satisfies the following two conditions:
  • $i$) No sample is taken when the channel is busy, , $S_{i+1}\geq D_{i}$, or equivalently,
    $S_{i+1}-Z_{i}=D_{i}\text{ with }Z_{i}\geq 0\text{ and }i\in\mathbb{N},$ (4)
    with $Z_{i}$ representing the sampling waiting time. From this assumption, we can also obtain that $D_{i}=S_{i}+Y_{i},\forall i\in\mathbb{N}$.
  • $ii$) The inter-sample times $G_{i}=S_{i+1}-S_{i}$ is a regenerative process [87, Section 6.1]: There is a sequence $0\leq g_{1}<g_{2}<\cdots$ of almost surely finite random integers such that the post-$g_{j}$ process $\{G_{{g_{j}}+i}\}_{i\in\mathbb{N}}$ has the same distribution as the post-$g_{1}$ process $\{G_{{g_{1}}+i}\}_{i\in\mathbb{N}}$ and is independent of the pre-$g_{j}$ process $\{G_{i}\}_{i\in\{0,1,\cdots g_{j}-1\}}$. Condition $ii$) implies that, almost surely [^4] This assumption also implies that the waiting time $Z_{i}$ is bounded, belonging to a subset of nature numbers with $Z_{i}\in\mathcal{Z}\subseteq\mathbb{N}$.
    $\lim\limits_{i\rightarrow\infty}S_{i}=\infty,~{}~{}~{}\lim\limits_{i\rightarrow\infty}D_{i}=\infty.$ (5)
In addition, we assume that the controlled policy satisfies the following condition: the action $a_{t}$ is updated only upon the delivery of a sample $X_{S_{i}}$ [^5] Throughout the time interval $t\in[D_{i},D_{i+1})$, the decision maker’s observations are fixed and consist only of $\{(X_{S_{j}},S_{j}):{j\leq i}\}$. Thus, we assume that the chosen actions remain constant in this period. , i.e.,
$a_{t}=A_{i},D_{i}\leq t\leq D_{i+1},i\in\mathbb{N},$ (6)
where $A_{i}\in\mathcal{A}$ is the updated controlled action upon the delivery of packet $(X_{S_{i}},S_{i})$. We consider a bounded cost function $\mathcal{C}(X_{t},a_{t})<\infty$, which represents the immediate cost incurred when action $a_{t}$ is taken in state $X_{t}$.
Under the above assumptions, the objective of the system is to design the optimal joint sampling and decision policies at each time slot, i.e., $\pi_{0},\pi_{1},\pi_{2}\cdots$, to minimize the long-term average cost, subject to a long-term average sampling frequency constraint:
Problem 1 (Joint Design of Sampling and Decision Processes under Sampling Frequency Constraint).
$\inf_{\pi_{0:\infty}}\limsup_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi_{0:\infty}}\left[\sum_{t=1}^{T}\mathcal{C}(X_{t},a_{t})\right]$ (7a)
[…]
where $\pi_{t}$ is the joint sampling and decision-making policy defined by (3), $f_{\text{max}}$ represents the maximum allowed sampling frequency, and the expectation $\mathbb{E}_{\pi_{0:\infty}}$ is taken over the stochastic processes $(X_{1},X_{2},\cdots)$ and $(Y_{0},Y_{1},\cdots)$ under given policies $\pi_{0:\infty}$.
In practice, the overhead for information updates will increase with the average sampling frequency. Hence, Problem 1 represents a tradeoff between remote decision-making utility and communication overhead. Since age $\Delta(k)$ is available at the decision maker as side information to facilitate more informed decision-making, we call this problem an age-aware remote MDP problem. This problem aims at determining the distribution of joint sampling and controlled actions $(u_{t},a_{t})$ based on the history $\mathcal{I}_{t}$, such that the long-term average cost subject to the sampling frequency constraint is minimized.
We remark that we will study Problem 1 both with (see Section IV) and without (see Section VI) the sampling frequency constraint (7b). To distinguish these two problems, we use $h^{\star}$ to denote the optimal value of the problem with the rate constraint and $\rho^{\star}$ to denote the optimal value of the problem without the sampling frequency constraint.
IV Optimal Sampling Without Rate Constraint: A Two-Layer Persample Solution
In this section, we address the unconstrained problem to determine $\rho^{\star}$. A series of theoretical results are developed through a divide-and-conquer approach. In subsection IV-A, we identify that the input to the policy $\pi_{t}$ is subject to the curse of history [88]. To address this, we analyze the sufficient statistics of the input to extract the most compressed representation that is necessary for decisions. In subsection IV-B, we rewrite the unconstrained problem into a non-linear fractional program. By utilizing a Dinkelbach-like method [89], we transform the non-linear fractional programming into an infinite horizon MDP problem given a Dinkelbach parameter. The problem then becomes finding the Dinkelbach parameter such that the optimal value of the standard MDP is zero, which is equivalently a root-finding problem. To search for the root, in subsection IV-C we review a typical two-layer nested algorithm, namely Bisec-RVI [1, Algorithm 2] [^6] The idea of the two-layer nested algorithm has been applied in [63, 20, 90] and [44] to achieve Age-optimal and Mean Square Error (MSE)-optimal sampling and scheduling.. We note that the inner-layer Relative Value Iteration (RVI) algorithm of Bisec-RVI may suffer from divergence and thus propose an improved Bisec-$\tau$-RVI algorithm to achieve provable convergence.
IV-A Sufficient Statistics of History
The policy $\pi_{t}$ maps the history $\mathcal{I}_{t}$ to the action simplex space $\mathcal{P}(\mathcal{A}\times\{0,1\})$. A significant challenge in solving $\pi_{t}$ is that the size of the history space $\mathcal{I}_{t}=\{(X_{k-\Delta(k)},\Delta(k),a_{k-1}^{S},a_{k-1}):{k\leq t}\}$ explodes exponentially as $t$ increases, known as the curse of history [88]. In this subsection, we analyze the sufficient statistic of $\mathcal{I}_{t}$ to address the curse of history. In decision theory, a sufficient statistic is defined as:
Definition 1.
A sufficient statistic of $\mathcal{I}_{t}$ is a function $\mathcal{H}_{t}(\mathcal{I}_{t})$, such that
$\min_{a_{t:T}}\mathbb{E}\left[\sum_{k=t}^{T}\mathcal{C}(X_{k},a_{k})|\mathcal{I}_{t}\right]=\min_{a_{t:T}}\mathbb{E}\left[\sum_{k=t}^{T}\mathcal{C}(X_{k},a_{k})|\mathcal{H}_{t}(\mathcal{I}_{t})\right]$ (8)
holds for any $T>t$.
This definition suggests that the decision making leveraging the sufficient statistics $\mathcal{H}_{t}(\mathcal{I}_{t})$ can achieve the same performance as using the complete history $\mathcal{I}_{t}$. Thus, the compression $\mathcal{H}_{t}(\mathcal{I}_{t})$ is sufficient for the agent to implement an optimal action, enabling the design of efficient policies that maintain optimality while overcoming the curse of history. The following Lemma establishes an efficient sufficient statistics of the history $\mathcal{I}_{t}$ in a piecewise manner.
Lemma 1.
(Sufficient Statistics). During the interval $t\in[D_{i},D_{i+1})$, $\mathcal{G}_{i}=(X_{S_{i}},Y_{i},A_{i-1})\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ is a sufficient statistic of $\mathcal{I}_{t}$. In addition, determining the optimal sampling actions $u_{t}$ under condition (4) is equivalent to determining the optimal sampling time $S_{i+1}$, or the optimal waiting time $Z_{i}$.
Proof.
The proof is detailed in the the short version of this work in [1, Appendix A]. ∎
The above Lemma 1 indicates that determining the optimal policy
$\pi_{t}:\mathcal{I}_{t}\rightarrow\mathcal{P}(\mathcal{A}\times\{0,1\})$ (9)
for Problem 1 is equivalent to solving for an alternative policy given as
$\phi_{t}:\mathcal{S}\times\mathcal{Y}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{A}\times\mathcal{Z}),\forall t\in[D_{i},D_{i+1}),i\in\mathbb{N},$ (10)
which is a mapping from the sufficient statistics $\mathcal{G}_{i}=(X_{S_{i}},Y_{i},A_{i-1})\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ to a distribution over the joint sampling and decision-making actions $(A_{i},Z_{i})\in\mathcal{A}\times\mathcal{Z}$. Because the Cartesian product $\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ is independent of time $t$, the exponential growth of $\mathcal{I}_{t}$ with respect to $t$ is thus mitigated. Consequently, the policy $\{\pi_{t}\}_{t=0}^{\infty}$ in Problem 1 can be replaced by an alternative policy $\{\phi_{t}\}_{t=0}^{\infty}$, which reformulates the unconstrained version of Problem 1 as:
Problem 2 (From History-Dependent to State-Dependent Policy).
$\rho^{\star}\triangleq\inf_{\boldsymbol{\psi}}\limsup_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\mathcal{C}(X_{t},a_{t})\right].$ (11)
Here, we use $\boldsymbol{\psi}$ to denote the policy $\{\phi_{t}\}_{t=0}^{\infty}$.
IV-B A Reformulation of Problem 2
As indicated in Lemma 1, the sufficient statistic of history is in a piecewise manner with respect to the time slot $t$. This allows us to shift our problem-solving perspective from a slot-by-slot analysis to an interval-by-interval approach. Assuming that the inter-sampling times $G_{i}$ constitute a regenerative process and given $\lim\limits_{i\rightarrow\infty}D_{i}=\infty$, we can reformulate the objective function in Problem 2 as:
$\limsup_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\mathcal{C}(X_{t},a_{t})\right]$ (12a)
$=\lim_{\it{D}_{\mathrm{n}}\rightarrow\infty}\frac{\mathbb{E}\left[\sum_{t=1}^{\it{D}_{\mathrm{n}}}\mathcal{C}(X_{t},a(t))\right]}{\mathbb{E}\left[\it{D}_{\mathrm{n}}\right]}$ (12b)
$=\lim_{\mathrm{n}\rightarrow\infty}\frac{\sum_{i=0}^{n-1}\mathbb{E}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{n-1}\mathbb{E}\left[\it{D}_{i+1}-\it{D}_{i}\right]}$ (12c)
$=\lim_{\mathrm{n}\rightarrow\infty}\frac{\sum_{i=0}^{n-1}\mathbb{E}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{n-1}\mathbb{E}\left[Y_{i+1}+Z_{i}\right]}.$ (12d)
We can then rewrite Problem 2 as the following interval-by-interval variant:
Problem 3 (Interval-by-Interval Reformulation).
$\rho^{\star}\triangleq\inf_{\boldsymbol{\psi}}\lim_{\mathrm{n}\rightarrow\infty}\frac{\sum_{i=0}^{n-1}\mathbb{E}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{n-1}\mathbb{E}\left[Y_{i+1}+Z_{i}\right]}.$ (13)
Problem 3 is a non-linear fractional program, which is challenging due to its fractional nature. To simplify this problem, we consider the following sequential decision process with Dinkelbach parameter $\lambda\geq 0$:
Problem 4 (Standard Infinite-Horizon Sequential Decision Process with Dinkelbach Parameter $\lambda$).
$U(\lambda)\triangleq$ (14)
$\inf_{\boldsymbol{\psi}^{\lambda}}\lim_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\left\{\mathbb{E}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]-\lambda\mathbb{E}\left[Z_{i}+Y_{i+1}\right]\right\}}.$
By similarly applying the Dinkelbach-like method for non-linear fractional programming [89], we can obtain the following lemma:
Lemma 2.
The following assertions hold:
(i). $\rho^{\star}\gtreqless\lambda\text{ if and only if }U(\lambda)\gtreqless 0\text{. }$
(ii). When $U(\lambda)=0$, the policy solutions to Problem 4 are equivalent to those of Problem 3.
(iii). $U(\lambda)=0$ has a unique root, with $U(\rho^{\star})=0$.
Proof.
The proof is detailed in the the short version of this work in [1, Appendix B]. ∎
This key lemma enables the formulation of the following two-layer nested approach to determine $\rho^{\star}$.
Figure 3: Bisection search to find the root of the implicit function $U(\lambda)$. The implicit function $U(\lambda)$ is approximated using a value iteration approach. The interval containing the root, denoted by $(\lambda_{\downarrow},\lambda_{\uparrow})$, is halved at each outer-layer iteration, and this process eventually converges to the unique root $\rho^{\star}$.
Input: Tolerence $\epsilon>0$, MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$
1 Initialization: $\lambda_{\uparrow}=\min_{a}\sum_{s\in\mathcal{S}}{\pi}_{a}(s)\cdot\mathcal{C}(s,a)$, $\lambda_{\downarrow}=\min_{s,a}\mathcal{C}(s,a)$;
2 while $\lambda_{\uparrow}-\lambda_{\downarrow}\geq\epsilon$ do
3       $\lambda=(\lambda_{\uparrow}+\lambda_{\downarrow})/2$;
4       Run RVI (Iteration 1) or $\tau$-RVI (Iteration 2) to solve $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ and calculate $U(\lambda)$;
5       if $U(\lambda)>0$ then
6             $\lambda_{\downarrow}=\lambda$;
7
8      else
9            $\lambda_{\uparrow}=\lambda$;
10
Output: $\rho^{\star}=\lambda$
Algorithm 1 Two-layer approaches for $\rho^{\star}$
IV-C Two-layer Approaches: Bisec-RVI and Bisec-$\tau$-RVI
Following Lemma 2, solving Problem 3 reduces to solving Problem 4 to determine the optimal value $U(\lambda)$ while simultaneously finding the unique root $\rho^{\star}$ of the implicit function $U(\lambda)$, which can be solved using a two-layer nested algorithm as shown in Fig. 3 and Algorithm 1. In the inner layer, value iteration approaches are applied to approximate the optimal value of Problem 4, $U(\lambda)$, by resolving the reformulated MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ detailed in subsection IV-C1. In the outer layer, a bisection search algorithm approximates the unique root $\rho^{\star}$. Conventionally, the RVI algorithm is applied in the inner layer to iteratively approximate $U(\lambda)$ [1, 20], but it proves to be divergent in our formulation (see Fig. 2). To address this limitation, we propose $\tau$-RVI in this paper and present a rigorous convergence analysis. The condition and the rationale of the convergence will be discussed in this section.
IV-C1 Inner-Layer MDP Given Dinkelbach Parameter $\lambda$
To solve the inner layer sequential decision process in Problem 4, we reformulate it as an equivalent standard infinite-horizon MDP. A standard MDP is typically described by a quadruple: the state space, the action space, the transition probability, and the cost function. This subsection details this quadruple. The MDP with Dinkelbach parameter $\lambda$ is denoted as $\mathscr{P}_{\mathrm{MDP}}(\lambda)$:
  • State Space: the state of the equivalent MDP is the sufficient statistics $\mathcal{G}_{i}=(X_{S_{i}},Y_{i},A_{i-1})\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$, as established in Lemma 1.
  • Action Space: the action space of the MDP is composed of the tuple $(Z_{i},A_{i})\in\mathcal{\mathcal{Z}\times\mathcal{A}}$, where $Z_{i}$ is the sampling waiting time and $A_{i}$ is the controlled action that controls the source.
  • Transition Probability: The transition probability is defined by $\Pr(\mathcal{G}_{i+1}|\mathcal{G}_{i},Z_{i},A_{i})$. We have the transition probability established in (15), whose detailed proof is given in the short version of this work [1, Appendix C]:
    $\Pr(\mathcal{G}_{i}=(s^{\prime},\delta^{\prime},a^{\prime})|\mathcal{G}_{i}=(s,\delta,a),Z_{i},A_{i})$ (15)
    $=\Pr(Y_{i+1}=\delta^{\prime})\cdot[\mathbf{P}_{a}^{\delta}\cdot\mathbf{P}_{A_{i}}^{Z_{i}}]_{s\times s^{\prime}}\cdot\mathbbm{1}\{a^{\prime}=A_{i}\},$
    where $[\mathbf{P}]_{s\times s^{\prime}}$ denotes the element located at the $s$-th row and $s^{\prime}$-th column of the matrix $\mathbf{P}$.
  • Cost Function: the cost function is typically a real-valued function over the state space and the action space. We denote it as $g(\mathcal{G}_{i},Z_{i},A_{i})$ and will show how to tailor the cost function to establish equivalence with Problem 4.
    Lemma 3.
    If the cost function is defined by
    $g(\mathcal{G}_{i},Z_{i},A_{i};\lambda)\triangleq q(\mathcal{G}_{i},Z_{i},A_{i})-\lambda f(Z_{i}),$ (16)
    where
    $q(\mathcal{G}_{i},Z_{i},A_{i})\triangleq$
    $\mathbb{E}\left[\sum_{s^{\prime}\in\mathcal{S}}\left[\sum_{t=0}^{Z_{i}+\mathrm{Y}_{i+1}-1}\mathbf{P}_{A_{i-1}}^{Y_{i}}\cdot\mathbf{P}_{A_{i}}^{t}\right]_{X_{S_{i}}\times s^{\prime}}\cdot\mathcal{C}(s^{\prime},A_{i})\right],$ (17)
    with the expectation $\mathbb{E}$ taken over the random delay $Y_{i+1}$, and
    $f(Z_{i})\triangleq\mathbb{E}\left[Y_{i+1}\right]+Z_{i}.$ (18)
    Then Problem $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ is equivalent to Problem 4.
    Proof.
    The proof is given in the short version of this work in [1, Appendix D]. ∎
In this paper, we establish that the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ preserves the critical unichain [^7] An MDP is unichain if the resulting Markov chain under any stationary policy is an aperiodic unichain. property under random communication delays and age-aware system configurations. While the primal MDP, defined by a finite state space $\mathcal{S}$, action space $\mathcal{A}$, and transition probabilities $\mathbf{P}_{a}$, is an aperiodic unichain, it remains unresolved whether this property transfers to $\mathscr{P}_{\text{MDP}}(\lambda)$. Direct verification presents significant computational overhead, as determining the unichain property is proven to be NP-hard [91]. To overcome this challenge, the following theorem demonstrates that the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ inherits the unichain property from the primal MDP, thereby guaranteeing the existence of an optimal stationary deterministic policy.
Theorem 1.
If the primal MDP, characterized by a finite state space $\mathcal{S}$, a finite action space $\mathcal{A}$, and the transition probability $\mathbf{P}_{a}$ is an aperiodic unichain, then the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ is unichain, and an optimal stationary deterministic policy [^8] A policy is deemed stationary when it is independent of the time slot $t$. A policy is deemed deterministic if it a deterministic mapping from the state space to the action space. When an optimal stationary policy exists, solving the policy $\boldsymbol{\psi}^{\lambda}$ is equivalent to solving simply a stationary policy $\phi^{\lambda}$. exists for $\mathscr{P}_{\text{MDP}}(\lambda)$.
Proof.
The proof is given in the short version of this work in [1, Appendix E]. ∎
For a unichain MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ in which an optimal stationary deterministic policy exists, one can establish the following Average Cost Optimality Equations (ACOE) [92, Eq. 4.1]:
$\mathbf{ACOE}:$ $V^{\star}(\gamma;\lambda)+U(\lambda)=\min_{A_{i},Z_{i}}\Big{\{}g(\gamma,Z_{i},A_{i};\lambda)+$
$\mathbb{E}[V^{\star}\left(\gamma^{\prime};\lambda\right)|\gamma,Z_{i},A_{i}]\Big{\}},\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$ (19)
where $V^{\star}(\gamma;\lambda)\in\mathbb{R}$ is the optimal value function and $U(\lambda)\in\mathbb{R}$ is the optimal long-term average value of $\mathscr{P}_{\text{MDP}}(\lambda)$ in (14). By solving $U(\lambda)$ and $V^{\star}(\gamma;\lambda)$ for $\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ from the ACOE (IV-C1), we can obtain the optimal stationary deterministic sampling and remote decision making policy for $\mathscr{P}_{\text{MDP}}(\lambda)$, defined as: $\phi^{*}_{\lambda}:\mathcal{S}\times\mathcal{Y}\times\mathcal{A}\rightarrow\mathcal{Z}\times\mathcal{A}$.
$\phi^{*}_{\lambda}(\gamma)=\mathop{\arg\min}_{Z_{i},A_{i}}$$\Big{\{}g(\gamma,Z_{i},A_{i};\lambda)+$
$\mathbb{E}[V^{\star}\left(\gamma^{\prime};\lambda\right)|\gamma,Z_{i},A_{i}]\Big{\}},\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$ (20)
The ACOE in (IV-C1) represents a series of nonlinear equations, which is mathematically intractable to solve explicitly. One can resort to the typical Dynamic Programming (DP)-like RVI algorithm [93] to iteratively generate the sequences $\{U_{K}(\lambda)\}^{K\in\mathbb{N}^{+}}$ and $\{V_{K}(\gamma;\lambda)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}^{K\in\mathbb{N}^{+}}$ that conditionally converge asymptotically to the solutions of the ACOE.
Iteration 1.
(RVI Algorithm [93]). For a given $\lambda$, the RVI starts with a given initial value $\{V_{0}(\gamma;\lambda)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}$ and computes $U_{K+1}(\lambda)$ and $V_{K+1}(\gamma;\lambda)$ iteratively:
$U_{K+1}(\lambda)$ $=\min_{A_{i},Z_{i}}\Big{\{}\,g(\gamma^{\text{r}},Z_{i},A_{i};\lambda)$
$\quad+\,\mathbb{E}\big{[}V_{K}(\gamma^{\prime};\lambda)\mid\gamma^{\text{r}},Z_{i},A_{i}\big{]}\Big{\}},$ (21a)
$V_{K+1}(\gamma;\lambda)$ $=\min_{A_{i},Z_{i}}\Big{\{}\,g(\gamma,Z_{i},A_{i};\lambda)$
$\quad+\,\mathbb{E}\big{[}V_{K}(\gamma^{\prime};\lambda)\mid\gamma,Z_{i},A_{i}\big{]}\Big{\}}$
$\quad-\,U_{K+1}(\lambda),\quad\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$ (21b)
where $\gamma^{r}\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ is an arbitrarily chosen reference state and the conditional expectation $\mathbb{E}$ is taken with respect to the conditional distribution of the next state $\gamma^{\prime}$ given the current state and the current action. The iterative process continues until a predefined convergence criterion is satisfied.
Figure 4: Algorithmic behavior of RVI versus $\tau$-RVI: Divergence mechanisms and comparative performance.
However, in our specific context, the convergence of the RVI algorithm is not necessarily guaranteed (see Example 1). To further investigate this issue and develop a convergent alternative, we present sufficient conditions for its convergence in the following Lemma.
Lemma 4.
(Restatement of [94, Proposition 4.3.2]). If an MDP satisfies the following conditions:
(a) the MDP is a unichain MDP;
(b) the optimal stationary policy for the MDP yields an aperiodic transition probability matrix;
then the sequences $\{U_{K}(\lambda)\}_{K\in\mathbb{N}}$ and $\{V_{K}(\gamma;\lambda)\}_{K\in\mathbb{N}}$ will converge to the solution to the ACOE (IV-C1):
$\lim_{K\rightarrow\infty}U_{K}(\lambda)=U(\lambda)$ (22)
$\lim_{K\rightarrow\infty}V_{K}(\gamma;\lambda)=V^{\star}(\gamma;\lambda),\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$
In the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$, condition (a) in Lemma 4 is ensured by Theorem 1. Nevertheless, the following counterexample demonstrates that condition (b) does not necessarily hold in $\mathscr{P}_{\text{MDP}}(\lambda)$. Specifically, even if the primal MDP characterized by $\left\langle\mathcal{S},\mathcal{A},\mathbf{P}_{a},\mathcal{C}\right\rangle$ is aperiodic, the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ may still exhibit periodicity, which will cause the RVI algorithm to diverge.
Example 1.
(A divergence example of RVI). Consider the parameter setup described in Appendix LABEL:appendixh where the delay is constant with $p=0$. In this case, the optimal policy for the transformed MDP $\mathscr{P}_{\text{MDP}}(10)$ is:
$\phi^{*}_{10}(\overbrace{0}^{X_{S_{i}}},\overbrace{10}^{Y_{i}},\overbrace{0}^{A_{i-1}})=(\overbrace{0}^{Z_{i}},\overbrace{1}^{A_{i}}),$ (23a)
$\phi^{*}_{10}(\overbrace{1}^{X_{S_{i}}},\overbrace{10}^{Y_{i}},\overbrace{0}^{A_{i-1}})=(\overbrace{0}^{Z_{i}},\overbrace{1}^{A_{i}}),$ (23b)
$\phi^{*}_{10}(\overbrace{0}^{X_{S_{i}}},\overbrace{10}^{Y_{i}},\overbrace{1}^{A_{i-1}})=(\overbrace{0}^{Z_{i}},\overbrace{0}^{A_{i}}),$ (23c)
$\phi^{*}_{10}(\overbrace{1}^{X_{S_{i}}},\overbrace{10}^{Y_{i}},\overbrace{1}^{A_{i-1}})=(\overbrace{0}^{Z_{i}},\overbrace{0}^{A_{i}}).$ (23d)
The optimal stationary policy in (23) induces the sub-state sequence $\{A_{i}\}_{i\in\mathbb{N}^{+}}$ to alternate in a $(0,1,0,1,\dots)$ ping-pong manner, as shown in Fig. 4-(a). This alternating behavior results in the RVI oscillations as shown in Fig. 4-(b).
Lemma 4 and Example 1 show that the RVI algorithm [93] may not asymptotically converge to $U(\lambda)$, and the reason is the periodic nature inherent in the transformed MDP $\mathscr{P}_{\text{MDP}}(\lambda)$. Consequently, the existing two-layer Bisec-RVI Algorithm (e.g., [1, Algorithm 1], [71, Algorithm 1], and [84, Algorithm 1]) cannot reliably determine the root $\rho^{\star}$. To circumvent this problem, we propose a new iterative approach, namely $\tau$-RVI, in Iteration 2. This approach eliminates the need for condition (b) in Lemma 4 but guarantees rigorous convergence.
Iteration 2.
($\tau$-RVI Algorithm). For a given $\lambda$ and a parameter $0<\tau\leq 1$, the $\tau$-RVI iteratively generate sequences $\{\tilde{U}_{K}(\lambda)\}^{K\in\mathbb{N}^{+}}$ and $\{\tilde{V}_{K}(\gamma;\lambda)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}^{K\in\mathbb{N}^{+}}$ with a starting initial value $\{\tilde{V}_{0}({\gamma;\lambda})\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}$.
$\tilde{U}_{K+1}(\lambda)$ $=\min_{A_{i},Z_{i}}\bigg{\{}\,g(\gamma^{\text{r}},Z_{i},A_{i};\lambda)$
$\quad\quad\quad+\tau\,\mathbb{E}\big{[}\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma^{\text{r}},Z_{i},A_{i}\big{]}\bigg{\}},$ (24a)
$\tilde{V}_{K+1}(\gamma;\lambda)$ $=(1-\tau)\,\tilde{V}_{K}(\gamma;\lambda)$
$\quad+\min_{A_{i},Z_{i}}\bigg{\{}\,g(\gamma,Z_{i},A_{i};\lambda)$
$\quad+\tau\,\mathbb{E}\big{[}\tilde{V}_{K}(\gamma^{\prime};\lambda)\mid\gamma,Z_{i},A_{i}\big{]}\bigg{\}}$
$\quad-\tilde{U}_{K+1}(\lambda),\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$ (24b)
where $\gamma^{\text{r}}{\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}$ is a predefined fixed reference state with initial condition $\tilde{V}_{0}(\gamma^{\text{r}};\lambda)=0$.
Remark 1.
If $\tau=1$, then $\tau$-RVI reduces to RVI.
Fig. 4-(b) illustrates the convergence of $\tau$-RVI across various values of the parameter $\tau$, compared to the standard RVI algorithm [93]. The results demonstrate that $\tau$-RVI overcomes oscillatory behavior encountered by RVI, with convergence rates depending on the selected $\tau$ values. A rigorous convergence analysis of $\tau$-RVI will be presented in in section IV-D.
IV-C2 Outer-Layer Bisection and Bounds on $\rho^{\star}$
In the outer layer of Algorithm 1, the search interval $(\lambda_{\downarrow},\lambda_{\uparrow})$ is bisected on a slow time scale to approximate the root of $U(\lambda)$. This process relies on the uniqueness of the root of $U(\lambda)$ established in Lemma 2. For the bisection search process, the complexity mainly depends on the initialization of the search interval $(\lambda_{\downarrow},\lambda_{\uparrow})$, which requires establishing upper and lower bounds on $\rho^{\star}$. To address this, we establish the bounds on $\rho^{\star}$ to initialize the bisection search:
Lemma 5 (Upper and Lower Bounds on $\rho^{\star}$).
The lower bound of $\rho^{\star}$ can be defined by the minimum value of the cost function, given by
$\rho^{\star}\geq\min_{s,a}\mathcal{C}(s,a).$ (25)
The upper bound of $\rho^{\star}$ can be defined by the minimum stationary cost achievable under a constant action, given by
$\rho^{\star}\leq\min_{a}\sum_{s\in\mathcal{S}}{\pi}_{a}(s)\cdot\mathcal{C}(s,a),$ (26)
where ${\pi}_{a}(s)$ represents the stationary distribution of state $s$, corresponding to the transition probability matrix $\mathbf{P}_{a}$.
Proof.
The proof is given in the short version of this work in [1, Appendix F ]. ∎
IV-D Convergence of $\tau$-RVI
In this subsection, we present convergence results for $\tau$-RVI in Iteration 2. We theoretically show that the generated sequences in $\tau$-RVI will asymptotically approach the solution to the ACOE (IV-C1). To quantify the convergence behavior, we define the relative error for the $K$-th iteration value $\tilde{U}_{K}(\lambda)$ with respect to the ACOE solution $U(\lambda)$ as:
$e_{\mathrm{U}}^{(K)}(\lambda)\triangleq|\tilde{U}_{K}(\lambda)-U(\lambda)|.$ (27)
Similarly, define the relative error for the $K$-th iteration value $\tilde{V}_{K}(\gamma;\lambda)$ with respect to the ACOE solution $V^{\star}(\gamma;\lambda)$ as:
$e_{\mathrm{V}}^{(K)}(\gamma;\tau,\lambda)\triangleq\left|\tilde{V}_{K}(\gamma;\lambda)-\frac{V^{\star}(\gamma;\lambda)}{\tau}\right|.$ (28)
The main convergence results for $\tau$-RVI are summarized in Theorem 2 below.
Theorem 2.
(Convergence of $\tau$-RVI). If the MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ is a unichain MDP, then the following limits hold true:
($i$). For $\forall\lambda\in\mathbb{R}$,
$\lim_{K\to\infty}e_{\mathrm{U}}^{(K)}(\lambda)=0.$ (29)
($ii$). For $\forall\lambda\in\mathbb{R}$, $\forall 0<\tau<1$, and $\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$,
$\lim_{K\to\infty}e_{\mathrm{V}}^{(K)}(\gamma;\tau,\lambda)=0.$ (30)
Proof Sketch.
The motivation behind (24) is to formulate an alternative MDP problem, denoted as $\widetilde{\mathscr{P}_{\text{MDP}}}(\lambda)$, by eliminating potential periodicity in the transition probabilities in (15). We denote the transition probability from state $i$ to state $j$, given action $a$, as $p_{ij}(a)$ in ${\mathscr{P}_{\text{MDP}}}(\lambda)$, and as $\widetilde{p_{ij}}(a)$ in $\widetilde{\mathscr{P}_{\text{MDP}}}(\lambda)$. The alternative transition probability $\widetilde{p_{ij}}(a)$ is defined as:
$\widetilde{p_{ij}}(a)\triangleq\begin{cases}\tau{p}_{ij}(a),&\text{if }i\neq j\\ 1-\tau+\tau p_{ii}(a),&\text{if }i=j.\end{cases}$ (31)
It is easy to verify that $\sum_{j}\widetilde{p_{ij}}(a)=1$ and $\widetilde{p_{ii}}(a)>0$ for $\forall i\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$, and thus the new MDP $\widetilde{\mathscr{P}_{\text{MDP}}}(\lambda)$ is aperiodic. Then, we can formulate the ACOE for the alternative $\widetilde{\mathscr{P}_{\text{MDP}}}(\lambda)$:
$\tilde{V}^{\star}(\gamma;\lambda)$ $+\tilde{U}(\lambda)=\min_{A_{i},Z_{i}}\Big{\{}g(\gamma,Z_{i},A_{i};\lambda)+$ (32)
$\sum_{\gamma^{\prime}}\widetilde{p_{\gamma\gamma^{\prime}}}(Z_{i},A_{i})\tilde{V}^{\star}(\gamma^{\prime};\lambda)\Big{\}},\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$
which can be solved by the traditional RVI with guaranteed convergence because of its unichain and aperiodic property. Then, comparing (IV-C1) and (32), we can establish the relationship that $\tau\tilde{V}^{\star}(\gamma;\lambda)={V}^{\star}(\gamma;\lambda)$ for $\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ and $\tilde{U}(\lambda)={U}(\lambda)$. See Appendix LABEL:proof:convergence:theorem2 for the detailed proof. ∎
The next theorem characterizes an upper bound on the relative error of the $\tau$-RVI Algorithm, whose proof is included in Appendix LABEL:proof:convergence:theorem2.
Theorem 3.
(Upper Bound of Relative Error). If the MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ is a unichain MDP, then up to iteration $K$, the relative error $e_{\mathrm{U}}^{(K)}(\lambda)$ is upper bounded above by
$e_{\mathrm{U}}^{(K)}(\lambda)\leq\frac{\tau M(1-\epsilon)^{(K-1)/L}}{1-(1-\epsilon)^{1/L}}=\mathcal{O}\left(\frac{1}{R^{K}}\right),$ (33)
where $M$ is a scaling factor, $L$ is defined by (LABEL:Ldefinition). The term $R=\frac{1}{(1-\epsilon)^{1/L}}>1$ captures the asymptotic convergence rate.
Theorem 3 demonstrates that the upper bound of the relative error decreases exponentially with respect to the number of iterations $K$. This indicates that the proposed method exhibits faster convergence, as the number of inner iterations required to achieve a given accuracy $\delta$ is at most logarithmic, i.e., $K\leq\mathcal{O}\left(\log(1/\delta)\right)$.
V Optimal Sampling Without Rate Constraint: A One-layer Primal-Dinkelbach Approach
The two-layer algorithm requires repeatedly executing the computation-intensive $\tau$-RVI to evaluate $U(\lambda)$ at the fast time scale and update $\lambda$ on the slow timescale. This results in high complexity since each update of $\lambda$ in the outer-layer search necessitates running the inner layer value iterations. To address this, we develop efficient one-layer iterations in this section that eliminate the need for outer-layer bisection search. The key idea here is to treat the constraint $U(\rho^{\star})=0$ as an intrinsic condition within the Markov Decision Process (MDP) $\mathscr{P}_{\mathrm{MDP}}(\rho^{\star})$, allowing us to directly establish the optimality equations, which take the form of fixed-point equations, as shown in [1]. In this work, we show through an example that the fixed-point operation in [1] may not converge due to the non-contractive nature of the operator (see Section V-A). To address divergence, we develop a new one-layer iterative algorithm that guarantees provable convergence to the fixed point (see Section V-B).
V-A Fixed-Point Equations and Iterations
We demonstrate here that the root finding process on $U(\lambda)$ is equivalent to solving the following non-linear equations, which we will then show to be fixed-point equations.
Theorem 4.
Solving Problem $\mathscr{P}_{\mathrm{MDP}}(\rho^{\star})$ with $U(\rho^{\star})=0$ is equivalent to solving the following nonlinear equations:
[…]
where $\gamma^{\mathrm{r}}\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ is the reference state and can be arbitrarily chosen.
Proof.
The proof is detailed in the short version of this work in [1, Appendix G]. ∎
In (34), there are $|\mathcal{S}|\times|\mathcal{Y}|\times|\mathcal{A}|+1$ variables, i.e., $\rho^{\star}$ and $W^{\star}(\gamma),\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$, which are matched by an equal number of equations. Solving non-linear equations is generally challenging, however, we establish that the equations (34) are fixed-point equations. Let $\mathbf{W}^{\star}$ denote the vector consisting of $W^{\star}(\gamma)$ for all $\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$, (34) can be succinctly represented as follows:
[…]
where $T(\cdot):\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}$ is a non-linear operator corresponding to the first equations of (35) and $H(\cdot):\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}\rightarrow\mathbb{R}^{+}$ is an operator corresponding to the second equation of (35). Substituting the second equation $\rho^{\star}=H(\mathbf{W}^{\star})$ into the first equation of (35) yields $\mathbf{W}^{\star}=T(\mathbf{W}^{\star},H(\mathbf{W}^{\star}))$. Define a new operator $G$ as $G(\mathbf{W})\triangleq T(\mathbf{W},H(\mathbf{W}))$, as implied by (35), we have the fixed-point equation:
$\mathbf{W}^{\star}=G(\mathbf{W}^{\star}).$ (36)
Then, as was done in [1, Algorithm 2], we can conduct the following fixed-point iterations to asymptotically approximate the solution to (34):
[…]
Typically, the convergence of fixed-point iterations is established using the Banach Contraction Mapping Theorem [86, Theorem 6]. According to this theorem, if the operator $Q(\cdot):\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}\rightarrow\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}$ is a contraction mapping, i.e., for any $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}$, there exists a constant $10<\alpha<$ 1 such that
$||G(\mathbf{W}_{1})-G(\mathbf{W}_{2})||_{2}\leq\alpha||\mathbf{W}_{1}-\mathbf{W}_{2}||_{2},$ (38)
then the operator $G$ has a unique fixed point, and the fixed-point iteration described in (37) converges to the solution $\mathbf{W}^{\star}$, where $\mathbf{W}^{\star}=G(\mathbf{W^{\star}})$.
However, the following counter-example shows that the operator $G:\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}\rightarrow\mathbb{R}^{|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|}$ may be a non-expansive mapping [^9] Non-expansive mapping indicates $\alpha=1$ in (38). Fig. 5 demonstrates that the fix-point operation in (37) may be non-expansive. rather than a contraction mapping, which results in potential divergence.
Example 2.
(Divergence Example of FPBI [1, Algorithm 2]). Consider the parameter setup described in Appendix LABEL:appendixh where the delay is constant with $p=0$. In this case, the fixed-point iteration starting from $\mathbf{W}^{0}=\mathbf{0}$ will diverge, as shown in Fig. 5.
Figure 5: Convergence comparison between FPBI [1, Algorithm 2] and OnePDSI (Iteration 3). FPBI becomes non-expansive and thus diverges. The proposed OnePDSI converges.
V-B Primal-Dinkelbach Synchronous Iteration (Proposed OnePDSI)
To overcome the divergence limitations of FPBI, we develop a novel one-layer iterative approach that guarantees rigorous convergence to the solution of the fixed-point equations (34). The details of the iteration are given as:
Iteration 3.
(OnePDSI): For a given $10<\kappa<$ 1, we can iteratively generate sequences $\{\rho_{K}\}^{K\in\mathbb{N}^{+}}$ and $\{\widetilde{W}_{K}(\gamma)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}^{K\in\mathbb{N}^{+}}$ with a starting initial value $\{\widetilde{W}_{0}(\gamma)\}_{\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}$:
$\rho_{K+1}$ $=\min_{A_{i},Z_{i}}\Bigg{\{}\frac{q(\gamma^{\text{r}},Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma^{\text{r}})\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}$
$+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma^{\text{r}},Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg{\}}+\widetilde{W}_{K}(\gamma^{\text{r}}),$ (39a)
$\widetilde{W}_{K+1}(\gamma)$ $=\min_{A_{i},Z_{i}}\Bigg{\{}\frac{q(\gamma,Z_{i},A_{i})-\kappa\widetilde{W}_{K}(\gamma)\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}$
$+\frac{\kappa\mathbb{E}\left[\widetilde{W}_{K}(\gamma^{\prime})\mid\gamma,Z_{i},A_{i}\right]\cdot\mathbb{E}[Y_{i}]}{f(Z_{i})}\Bigg{\}}$
$+\widetilde{W}_{K}(\gamma)-\rho_{K+1},\quad\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$ (39b)
where $\gamma^{\text{r}}{\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}}$ is a fixed reference state with an initial condition $\widetilde{W}_{0}(\gamma^{\text{r}})=0$.
In what follows, we present a comprehensive convergence analysis demonstrating that OnePDSI ensures robust and provable convergence.
V-C Convergence of OnePDSI
In this subsection, we theoretically demonstrate that the sequences in OnePDSI will approach the solution to (34). To quantify the convergence, we define the relative error of ${\rho_{K}}$ at iteration $K$ as:
$e_{\rho}^{(K)}=|\rho_{K}-\rho^{\star}|.$ (40)
Meanwhile, define the relative error of the sequence $\widetilde{W}_{K}(\gamma)$ at $K$-th iteration as
$e_{\mathrm{W}}^{(K)}(\gamma;\kappa)=\left|\widetilde{W}_{K}(\gamma)-\frac{{W}^{\star}(\gamma)}{\kappa\cdotp\mathbb{E}[Y_{i}]}\right|.$ (41)
The following theorem demonstrates the convergence of OnePDSI.
Theorem 5.
(Convergence of OnePDSI). If the transformed MDP is unichain and $10<\kappa<$ 1, then the OnePDSI in (39) is convergent, with:
$\lim_{k\to\infty}e_{\rho}^{(K)}=0,$ (42)
$\lim_{k\to\infty}e_{\mathrm{W}}^{(K)}(\gamma)=0,\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}.$
Proof.
See Appendix LABEL:appendixJ. ∎
The next theorem characterizes an upper bound on the relative error of the proposed OnePDSI, whose proof is provided in Appendix LABEL:appendixJ:
Theorem 6.
(Upper Bound of Relative Error). If the MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ is a unichain MDP, then up to iteration $K$, the relative error $e_{\mathrm{\rho}}^{(K)}(\gamma)$ is upper bounded above by
$e_{\mathrm{U}}^{(K)}(\lambda)\leq\frac{\tau M(1-\epsilon)^{(K-1)/L}}{1-(1-\epsilon)^{1/L}}=\mathcal{O}\left(\frac{1}{R^{K}}\right),$ (43)
where $M$ is a scaling factor, $L$ is defined by (LABEL:181). The term $R=\frac{1}{(1-\epsilon)^{1/L}}$ captures the asymptotic convergence rate.
Theorem 6 demonstrates that the upper bound of the relative error decreases exponentially with respect to the number of iterations $K$. This indicates that the number of inner iterations required to achieve a given optimality gap $\delta$ is at most logarithmic, i.e., $K\leq\mathcal{O}\left(\log(1/\delta)\right)$.
VI Optimal Sampling With Rate Constraint: A Typical Three-Layer Approach
In this section, we investigate the optimal sampling and decision-making policy that minimizes the long-term average cost under a sampling frequency constraint, as formulated in Problem 1. Our goal is to derive this optimal policy and its corresponding value $h^{\star}$.
VI-A Lagrangian Dual Techniques
Following the steps in (11)-(14) and applying Dinkelbach’s method for non-linear fractional programming as in [89] and [20, Lemma 2], the problem of determining the optimal policy for Problem 1 is equivalent to solving the following alternative problem given the Dinkelbach parameter $\lambda$:
Problem 5 (Standard Infinite-Horizon Constrained Markov Decision Process (CMDP) with Dinkelbach Parameter $\lambda$).
$H(\lambda;f_{\max})\triangleq$ (44)
$\inf_{\boldsymbol{\psi}}\lim_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\left\{\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]-\lambda\mathbb{E}_{\boldsymbol{\psi}}\left[Z_{i}+Y_{i+1}\right]\right\}}$
[…]
The following lemma characterizes the relationship between $h^{\star}$ and the optimal value of Problem 5. The proof follows directly from modifying the policies in [1, Appendix B] to satisfy the sampling frequency constraint, and is therefore omitted.
Lemma 6.
The following assertions hold:
(i). $h^{\star}\gtreqless\lambda\text{ if and only if }H(\lambda;f_{\max})\gtreqless 0\text{. }$
(ii). When $H(\lambda;f_{\max})=0$, the solution to Problem 5 coincide with those of Problem Problem 1.
(iii). $H(\lambda;f_{\max})=0$ has a unique root, and the root is $h^{\star}$.
With Lemma 6 in hand, solving Problem 1 is equivalent to solving for the root of the implicit function $H(\lambda;f_{\max})$. To obtain $H(\lambda;f_{\max})$ given $\lambda$, we first transform the CMDP into an unconstrained Lagrangian MDP problem. Specifically, define the Lagrange Dual Function as:
$\mathcal{L}(\boldsymbol{\psi};\theta,\lambda,f_{\max})=\frac{\theta}{f_{\max}}+$ (45)
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\boldsymbol{\psi}}\left\{\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-(\lambda+\theta)\left(Z_{i}+Y_{i+1}\right)\right\},$
where $\theta\geq 0$ is the Lagrangian multiplier. Let the Lagrangian Dual Function defined as:
$\Upsilon(\theta,\lambda;f_{\max})\triangleq\inf_{\boldsymbol{\psi}}\mathcal{L}(\boldsymbol{\psi};\theta,\lambda,f_{\max}).$ (46)
Since an optimal stationary deterministic policy exists for the MDP problem $\inf_{\boldsymbol{\psi}}\mathcal{L}(\boldsymbol{\psi};\theta,\lambda,f_{\max})$ (as indicated in Theorem 1), we can use a short-hand notation $\phi$ to denote $\boldsymbol{\psi}$, and the Lagrangian Dual Problem of Problem 5 is:
Problem 6 (Lagrangian Dual Problem).
$d(\lambda;f_{\max})=\max_{\theta\geq 0}\Upsilon(\theta,\lambda;f_{\max}),$ (47)
where $\Upsilon(\theta,\lambda;f_{\max})=\inf_{\phi}\mathcal{L}(\phi;\theta,\lambda,f_{\max})$
The weak duality principle [95, Chapter 5.2.2] implies that $d(\lambda;f_{\max})$ is a lower bound of $H(\lambda;f_{\max})$, i.e., $d(\lambda;f_{\max})\leq H(\lambda;f_{\max})$. In the following lemma, we establish the conditions where the strong duality holds true and thus $d(\lambda;f_{\max})=H(\lambda;f_{\max})$. Under these conditions, it is sufficient to solve $d(\lambda;f_{\max})$ to obtain $H(\lambda;f_{\max})$.
Lemma 7.
(Restatement of [95, Chapter 5.5.3]) The duality gap between Problem 5 and Problem 6 is zero, i.e., $d(\lambda;f_{\max})=H(\lambda;f_{\max})$, if and only if for any given $\lambda$, we can find $\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}$ and $\theta^{\star}_{\lambda}$ such that the Karush–Kuhn–Tucker (KKT) conditions are satisfied:
$\theta^{\star}_{\lambda}\geq 0,$ (48a)
$\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}=\mathop{\arg\min}_{\phi}\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=0}^{n-1}\mathbb{E}_{\phi}\Bigg{\{}\sum_{t=D_{i}}^{D_{i+1}-1}\mathcal{C}\left(X_{t},A_{i}\right)-$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad(\lambda+\theta^{\star}_{\lambda})\left(Z_{i}+Y_{i+1}\right)\Bigg{\}}+\frac{\theta^{\star}_{\lambda}}{f_{\max}},$ (48b)
$\lim_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\mathbb{E}_{\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}}\left[Z_{i}+Y_{i+1}\right]}\geq\frac{1}{f_{\max}},$ (48c)
$\theta^{\star}_{\lambda}\left\{\lim_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\mathbb{E}_{\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}}\left[Z_{i}+Y_{i+1}\right]}-\frac{1}{f_{\max}}\right\}=0.$ (48d)
By leveraging Lemma 7, we reformulate the constrained problem as an unconstrained Problem 6. Given a fixed Dinkelbach parameter $\lambda$, the goal of Problem 6 is to determine the saddle point $(\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}},\theta^{\star}_{\lambda})$ of the function $\mathcal{L}(\phi;\theta,\lambda)$. The inner layer of Problem 6 is a standard MDP problem with fixed Lagrangian multiplier $\theta\geq 0$ and Dinkelbach parameter $\lambda$, while the outer layer seeks the optimal Lagrangian multiplier $\theta^{\star}_{\lambda}$ that maximizes the Lagrangian Dual Function (46) under the KKT conditions (48a)-(48d).
VI-B Three-Layer Solutions and the Structure of Optimal Policies
In this subsection, we propose a three-layer algorithm outlined in Algorithm 2. The basic framework of this algorithm is inspired by [85, Section IV.C]. This algorithm consists of inner, middle, and outer layers, whose implementations are detailed below.
VI-B1 Inner-Layer: A Standard MDP Given $\lambda$ and $\theta$
For any given $\theta$ and $\lambda$, the inner layer $\inf_{\phi}\mathcal{L}(\phi;\theta,\lambda,f_{\max})$ is a standard unconstrained infinite horizon MDP as defined in IV-C1, which is denoted by $\mathscr{P}_{\text{MDP}}(\theta+\lambda)$ with the optimal value $U(\lambda+\theta)$. This problem can be efficiently solved using the $\tau$-RVI algorithm, as outlined in Iteration 2. Notably, [85, Section IV.C] applies a value iteration process in the inner layer, while we apply the $\tau$-RVI to ensure the rigorous convergence in our context.
VI-B2 Middle-Layer: Update Lagrangian multiplier $\theta$ Given $\lambda$
Given a Dinkelbach parameter $\lambda$, the middle layer involves solving $\max_{\theta\geq 0}\Upsilon(\theta,\lambda;f_{\max})$ in Problem 6 to accurately approximate $d(\lambda;f_{\max})$, which searches for the optimal $\theta^{\star}_{\lambda}$ and its corresponding optimal policy $\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}$ that satisfy the KKT conditions in Lemma 7. Different from [85, Section IV.C] where the sub-gradient method is employed to update the Lagrangian multiplier leveraging the convexity of the Lagrangian Dual function, we conduct a monotonicity analysis and explicitly state the conditions for the optimal multiplier. We begin by introducing the following notations that help clarify the role of the multiplier $\theta$ given a fixed $\lambda$. Specifically, denote
$\mathcal{Q}^{\lambda+\theta}\triangleq\limsup_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\mathbb{E}_{\phi^{\star}_{\lambda+\theta}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]},$ (49)
$\mathcal{F}^{\lambda+\theta}\triangleq\liminf_{\mathrm{n}\rightarrow\infty}\frac{1}{n}{\sum_{i=0}^{n-1}\mathbb{E}_{\phi^{\star}_{\lambda+\theta}}\left[Z_{i}+Y_{i+1}\right]},$ (50)
where $\phi^{\star}_{\lambda+\theta}$ is the optimal stationary deterministic policy for the unconstrained problem $\mathscr{P}_{\text{MDP}}(\lambda+\theta)$ given in IV-C1. The subsequent Lemma presents key properties of $\mathcal{F}^{\lambda+\theta}$, $\mathcal{Q}^{\lambda+\theta}$, $U(\lambda+\theta)$, and $\Upsilon(\theta,\lambda;f_{\max})$ with respect to $\theta$, which are useful for the explicit solutions.
Lemma 8.
(A variant of [96, Lemma 3.1]) The following assertions hold true:
($i$) $U({\lambda+\theta})$ is non-increasing with respect to $\theta$;
($ii$) $\mathcal{Q}^{\lambda+\theta}$ and $\mathcal{F}^{\lambda+\theta}$ are non-decreasing functions with respect to $\theta$;
($iii$) $\mathcal{Q}^{\lambda+\theta}$ and $\mathcal{F}^{\lambda+\theta}$ are both step functions with respect to $\theta$;
($iv$) If $\mathcal{F}^{\lambda+\theta}\geq 1/f_{\max}$, $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with respect to $\theta$.
Proof.
Part ($i$) and part ($ii$) are supported by [96, Lemma 3.1]. Part ($iii$) is supported by [96, Lemma 3.2]. See Appendix LABEL:appendixkv3 for the detailed proof of part ($iv$). ∎
Lemma 8 leads to the following corollary, which provides an explicit solution for the optimal value $\theta_{\lambda}^{\star}$ in Problem 6 that satisfies the KKT conditions (48a)-(48d).
Figure 6: Illustrations of Case ($i$) and case ($ii$) in Corollary 1. In this figure, $\mathcal{F}^{\lambda+\theta}$ is a non-increasing step function with respect to $\theta$, as Lemma 8-($ii$) and Lemma 8-($iii$) indicate. In addition, $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with respect to $\theta$ if $\mathcal{F^{\lambda+\theta}}\geq 1/f_{\max}$, as Lemma 8-($iv$) indicates.
Corollary 1.
Denote $\mathcal{F}^{\lambda^{+}}$ as the right limit [^10] Since $\mathcal{F}^{\lambda}$ is a step function, it does not necessarily follow that $\mathcal{F}^{\lambda^{+}}=\mathcal{F}^{\lambda}$. Specifically, when $\lambda$ is a break point, we have $\mathcal{F}^{\lambda^{+}}>\mathcal{F}^{\lambda}$. of $\mathcal{F}^{\lambda}$:
$\mathcal{F}^{\lambda^{+}}=\lim_{\Delta\lambda\rightarrow 0}\mathcal{F}^{\lambda+\Delta\lambda},$ (51)
The following assertions hold true:
$(i)$. If $\mathcal{F}^{\lambda^{+}}\geq 1/f_{\max}$, then $\theta_{\lambda}^{\star}=0$;
$(ii)$. If $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$, then $\theta_{\lambda}^{\star}$ is equal to a positive break point $\theta^{\text{tr}}>0$, which satisfies:
$\mathcal{F}^{(\theta^{\text{tr}}+\lambda)^{-}}<\frac{1}{f_{\max}}\leq\mathcal{F}^{(\theta^{\text{tr}}+\lambda)^{+}},$ (52)
Proof Sketch.
If $\mathcal{F}^{\lambda+}\geq\frac{1}{f_{\max}}$, which is illustrated in the right panel of Fig. 6, the function $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing with $\theta$ for $\theta\geq 0$. In this case, the maximum value of $\Upsilon(\theta,\lambda;f_{\max})$ is obtained at $\theta_{\lambda}^{\star}=0$. If $\mathcal{F}^{\lambda+}<\frac{1}{f_{\max}}$, the feasible region under the KKT condition (48c) is shown in the left panel of Fig. 6 with $\theta\geq\theta^{\text{tr}}$. In this feasible region, the function $\Upsilon(\theta,\lambda;f_{\max})$ is non-increasing in $\theta$. Therefore, the maximum value of $\Upsilon(\theta,\lambda;f_{\max})$ is obtained at $\theta_{\lambda}^{\star}=\theta^{\text{tr}}$. A detailed proof is provided in Appendix LABEL:appendixlv2. ∎
Having established Corollary 1, our remaining task is to identify the threshold value $\theta^{\text{tr}}$ that satisfies (52) under the condition that $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$. Given that $\mathcal{F}^{\lambda+\theta}$ is non-decreasing in $\theta$, we can apply a bisection search to gradually converge on the threshold $\theta^{\text{tr}}$. Once this threshold value is obtained, setting $\theta_{\lambda}^{\star}=\theta^{\text{tr}}$ ensures that the KKT conditions (48a)–(48c) are satisfied. What remains is to determine the optimal policy $\phi_{\lambda+\theta^{\star}_{\lambda}}$ that guarantees the KKT condition (48d). The structure of this optimal policy is provided in the following theorem.
Theorem 7.
(Structure of the optimal policy) The following assertions hold true:
($i$). If $\mathcal{F}^{\lambda^{+}}\geq 1/f_{\max}$, the optimal policy is a stationary deterministic policy $\phi^{\star}_{\lambda}$ determined in (IV-C1);
($ii$). If $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$ and $\mathcal{F}^{(\lambda+\theta^{\star}_{\lambda})^{+}}=1/f_{\max}$, the optimal policy is a stationary deterministic policy $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})}$;
($iii$). If $\mathcal{F}^{\lambda^{+}}<1/f_{\max}$ and $\mathcal{F}^{(\lambda+\theta^{\star}_{\lambda})^{+}}>1/f_{\max}$, the optimal policy is a random mixture of two stationary deterministic policy $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})^{+}}$ and $\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})^{-}}$:
$\phi^{\star}_{\lambda+\theta^{\star}_{\lambda}}(\gamma)=\begin{cases}\phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})^{+}}(\gamma),&\text{w.p. }\eta\\ \phi^{\star}_{(\lambda+\theta^{\star}_{\lambda})^{-}}(\gamma),&\text{w.p. }1-\eta\end{cases},\text{ for }\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$ (53)
where $\eta$ is a randomization factor given by
$\eta=\frac{\frac{1}{f_{\max}}-\mathcal{F}^{\lambda+\theta^{\star}_{\lambda}}}{\mathcal{F}^{(\lambda+\theta^{\star}_{\lambda})^{+}}-\mathcal{F}^{\lambda+\theta^{\star}_{\lambda}}}.$ (54)
Proof.
See Appendix LABEL:appendixkM. ∎
With Theorem 7, we can determine the optimal policy once the optimal $\theta_{\lambda}^{\star}$ is obtained. In this way, the optimal value $d(\lambda;f_{\max})$ is determined as:
$d(\lambda;f_{\max})$ $=\Upsilon(\theta_{\lambda}^{\star},\lambda;f_{\max})$ (55)
=$=\begin{cases}U(\lambda)&\text{if }\mathcal{F}^{\lambda^{+}}\geq\frac{1}{f_{\max}}\\ U(\lambda+\theta_{\lambda}^{\star})+\frac{\theta_{\lambda}^{\star}}{f_{\max}}&\text{if }\mathcal{F}^{\lambda^{+}}<\frac{1}{f_{\max}}\end{cases}.$
The following subsection aims at searching the root of $d(\lambda;f_{\max})$.
Input: Tolerence $\epsilon_{1},\epsilon_{2}>0$, MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$, maximum sampling frequency $f_{\max}$
1 Initialization: sufficiently large $\lambda_{\uparrow}$, and $\lambda_{\downarrow}=\min_{s,a}\mathcal{C}(s,a)$;
// Outer-Layer Bisection Search on Dinkelbach parameter $\lambda$
2 while $\lambda_{\uparrow}-\lambda_{\downarrow}\geq\epsilon_{1}$ do
3      $\lambda=(\lambda_{\uparrow}+\lambda_{\downarrow})/2$;
4       Run $\tau$-RVI to solve $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ and calculate $\mathcal{F}^{\lambda^{+}}$ and $U(\lambda)$;
5       if $\mathcal{F}^{\lambda^{+}}\geq 1/f_{\max}$ then
6            $d(\lambda;f_{\max})=U(\lambda)$;
7      else
8             Initialization: sufficiently large $\theta_{\uparrow}$, and $\theta_{\downarrow}=0$;
            // Middle-Layer Bisection Search on Lagrangian multiplier $\theta$
9             while $\theta_{\uparrow}-\theta_{\downarrow}\geq\epsilon_{2}$ do
10                   $\theta=(\theta_{\uparrow}+\theta_{\downarrow})/2$;
                  // Inner-Layer MDP Given $\lambda$ and $\theta$
11                  Run $\tau$-RVI to solve $\mathscr{P}_{\text{MDP}}(\lambda+\theta)$ and calculate $\mathcal{F}^{(\lambda+\theta)^{+}}$ and $U(\lambda+\theta)$;
12                   if $\mathcal{F}^{(\lambda+\theta)^{+}}\geq 1/f_{\max}$ then
13                         $\theta_{\uparrow}=\theta$;
14
15                  else
16                        $\theta_{\downarrow}=\theta$;
17
18            $d(\lambda;f_{\max})=U(\lambda+\theta)+\frac{\theta}{f_{\max}}$;
19
20      if $d(\lambda;f_{\max})>0$ then
21             $\lambda_{\uparrow}=\lambda$;
22
23      else
24            $\lambda_{\downarrow}=\lambda$;
25
Output: $h^{\star}=\lambda$
Algorithm 2 A Three-layer Algorithm (A Variant of [85, Section IV.C])
VI-B3 Outer-Layer: Update Dinkelbach parameter $\lambda$
With the approximation of $d(\lambda;f_{\max})$ in hand, the outer layer updates $\lambda$ in a bisection-search fashion by leveraging Lemma 6 to finally approach the root $h^{\star}$ such that $d(h^{\star};f_{\max})=0$. The flow of the algorithm is demonstrated in Algorithm 2.
VII Optimal Sampling With Rate Constraint: One-Layer Iteration is All You Need
Input: MDP $\mathscr{P}_{\mathrm{MDP}}(\lambda)$ and maximum sampling frequency $f_{\max}$
1 Run OnePDSI in Iteration 3 to obtain $\rho^{\star}$;
2 Calculate the left limit $\mathcal{F}^{(\rho^{\star})^{-}}$;
3 if $\mathcal{F}^{(\rho^{\star})^{-}}\geq 1/f_{\max}$ then
      $h^{\star}=\rho^{\star}$;
        // Case ($i$) of Theorem 8
4
5else
6      Solve LP Problem 7 to obtain $Q^{\star}(f_{\max})$;
       Calculate the root $h^{\star}=f_{\max}\cdot Q^{\star}(f_{\max})$;
        // Case ($ii$) of Theorem 8
7
Output: $h^{\star}$
Algorithm 3 Proposed One-layer QuickBLP
In the three-layer algorithm, each update of $\lambda$ or $\theta$ necessitates solving the inner-layer MDP using the $\tau$-RVI algorithm. This process incurs a high computational complexity due to the repeated execution of the $\tau$-RVI algorithm required to iteratively search and optimize the parameters $\lambda$ and $\theta$.
VII-A QuickBLP
In this section, we design a one-layer two-stage hierarchical algorithm namely, QuickBLP, which reduces the computational complexity by explicitly solving for the root $h^{\star}$. Rather than relying on iterative bisection search to find the root $h^{\star}$, our approach leverages a direct structural exploration of the solution space, allowing us to bypass the need for multiple $\tau$-RVI executions. Consequently, the QuickBLP solves this problem in two stages. The first stage solves a Bellman variant and the second stage explicitly expresses the root $h^{\star}$ as a function of the solution to an LP problem, thus directly obtaining the root $h^{\star}$ in a more computationally efficient manner. The main structural results are summarized in the following theorem.
Theorem 8.
(Structural Results of the root $h^{\star}$) The following assertions hold true:
($i$). If the root of $U(\rho)$, denoted as $\rho^{\star}$, satisfies $\mathcal{F}^{(\rho^{\star})^{-}}\geq 1/f_{\max}$, then the optimal value of Problem 1 is $h^{\star}=\rho^{\star}$, and the optimal policy for Problem 1 is $\phi_{\rho^{\star}}^{\star}$, as defined in (IV-C1);
($ii$). If $\mathcal{F}^{(\rho^{\star})^{-}}<1/f_{\max}$, the optimal value of Problem 1 is
$h^{\star}=f_{\max}\cdot Q^{\star}(f_{\max}),$ (56)
where $Q^{\star}(f_{\max})$ is the optimal value of the following Linear Programming:
Problem 7 (Linear Programming Reformulation).
$Q^{\star}(f_{\max})=\min_{\mathbf{x}}\sum_{\gamma,z,a}q(\gamma,z,a)x(\gamma,z,a)$ (57a)
$\mathrm{s.t.}\quad$ $\sum_{\gamma,z,a}f(z)x(\gamma,z,a)=\frac{1}{f_{\max}},$ (57b)
$\sum_{z,a}x(\gamma^{\prime},z,a)=\sum_{\gamma,z,a}p(\gamma^{\prime}|\gamma,z,a)x(\gamma,z,a),$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\forall\gamma^{\prime}\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},$ (57c)
$\sum_{\gamma,z,a}x(\gamma,z,a)=1,$ (57d)
$x(\gamma,z,a)\geq 0,~{}~{}\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},z\in\mathcal{Z},a\in\mathcal{A},$ (57e)
and the corresponding optimal policy is a randomized policy given by:
$\phi^{\star}(\gamma)=(z,a),\text{w.p. }\frac{x(\gamma,z,a)}{\sum_{\zeta,\alpha}x(\gamma,\zeta,\alpha)},$ (58)
$\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},z\in\mathcal{Z},a\in\mathcal{A}.$
Proof.
See Appendix LABEL:appendixkv2. ∎
Theorem 8 leads to a one-layer two-stage algorithm presented in Algorithm 3. In the first stage, the algorithm solves the unconstrained MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ and determines the root of $U(\lambda)$, denoted by $\rho^{\star}$. This root is obtained by implementing OnePDSI in Iteration 3. If the condition specified in part ($i$) of Theorem 8 holds true, then the root $h^{\star}$ is immediately found as $h^{\star}=\rho^{\star}$. If the condition of part ($i$) of Theorem 8 is not met, the algorithm proceeds to the second stage. Here, the LP problem [^11] Once an LP problem is established, we can say the solution is achieved. This is based on the mature techniques such as the simplex algorithm and the interior point method for solving LP problems. Problem 7 is solved. The solution of this LP provides the optimal value $Q^{\star}(f_{\max})$. Finally, the root $h^{\star}$ is explicitly determined as $h^{\star}=f_{\max}\cdot Q^{\star}(f_{\max})$. Compared to the previous Algorithm 2, this newly proposed algorithm eliminates the need for multiple executions of the $\tau$-RVI for searching $h^{\star}$ and $\theta^{\star}_{h^{\star}}$. Table II compares the two algorithms.
VII-B Sensitivity Analysis and Sampling Frequency Threshold
To better understand the relationship between $h^{\star}$ and the maximum rate constraint parameter $f_{\max}$, we conduct a sensitivity analysis in this subsection to explore how $h^{\star}$ depends on the maximum rate constraint with parameter $f_{\max}$. The key result is stated in the following Lemma, whose proof can be found in Appendix LABEL:apposa.
Theorem 9.
(Sensitivity Analysis). Define $f_{\max}^{T}$ as:
$f_{\max}^{T}\triangleq\frac{1}{\mathcal{F}^{(\rho^{\star})^{-}}},$ (59)
the following assertions hold true:
($i$). If $f_{\max}\geq f_{\max}^{T}$, $h^{\star}$ is independent of $f_{\max}$;
($ii$). If $f_{\max}<f_{\max}^{T}$, $h^{\star}=f_{\max}Q^{\star}(f_{\max})$ is monotonically non-increasing with $f_{\max}$, with the derivative given as:
$\frac{\mathrm{d}h^{\star}}{\mathrm{d}f_{\max}}=U(\lambda^{\star}),$ (60)
where
[…]
Furthermore, the condition that distinguishes the two cases in Lemma 9 directly establishes the following corollary, which characterizes the phenomenon where additional sampling does not contribute to further decision-making performance.
Corollary 2.
(Sampling frequency threshold) When the sampling frequency $f_{\max}$ exceeds or equals the threshold $f_{\max}^{T}$, i.e., $f_{\max}\geq f_{\max}^{T}$, any further increase in sampling frequency provides no improvement in decision performance.
TABLE II: Comparing Algorithm 2 and Algorithm 3
Aspect Algorithm 2 [85, Section IV.C] Algorithm 3 (Proposed QuickBLP)
Loop Structure Three nested layers Single-layer with two stages
Number of Iterations $\log\frac{1}{\epsilon_{1}}\cdot\log\frac{1}{\epsilon_{2}}$ $\tau$-RVI calls 1 OnePDSI call + 1 optional LP call
(LP is triggered only if $\mathcal{F}^{(\rho^{\star})^{-}}<1/f_{\max}$)
Parameter Sensitivity Depends on tolerance parameter $\epsilon_{1}$, $\epsilon_{2}$ Parameter-free (Solves $h^{\star}$ directly)
Note: Both $\tau$-RVI and OnePDSI have exponential convergence rates where the optimality gap decreases exponentially in the worst case.
VIII Simulation Results
VIII-A Comparing Benchmark
VIII-A1 Sampling Policies
In this paper, the sampling policy that is co-designed with the goal-oriented remote decision making to minimize the long-term average cost in Problem 1 is referred to as “Goal-oriented sampling”. We compare this sampling policy with the following benchmarks:
  • Uniform sampling: The sampling is activated periodically with a constant sampling interval $d\in\mathbb{N}^{+}$. In this case, the sampling process follows $S_{i+1}-S_{i}=d$ for $\forall i$ (see [4, Section VI] for a detailed system description of uniform sampling with queuing and random service delay). Since the sampling interval $d$ is limited to integer values in the discrete-time MDP setting, the corresponding sampling frequency $1/d$ also takes discrete values. Hence, the simulation yields a performance curve comprising discrete points, as shown in Fig. 8.
  • Zero-wait sampling: An update is transmitted once the previous update is delivered, i.e., $Z_{i}=0$ for $\forall i$. This policy achieves the maximum throughput. The sampling frequency for zero-wait sampling is feasible only if $f_{\operatorname{max}}\geq 1/\mathbb{E}[Y_{i}]$.
  • Constant-wait sampling: Waiting before transmitting is reported to be a good alternative for information freshness [97]. Here we consider $Z_{i}=z$, where $z\in\mathbb{N}^{+}$ is a constant waiting time. The sampling frequency for constant-wait sampling is $1/(z+\mathbb{E}[Y_{i}])$, which imposes a feasibility condition: $f_{\max}\geq 1/(z+\mathbb{E}[Y_{i}])$.
  • AoI-optimal sampling: The AoI-optimal policy determines $Z_{i}$ by [63, Theorem 4], which is a threshold-based policy $Z_{i}=\max(0,\beta-Y_{i})$, where $\beta$ is the solution to the following equations:
    $\mathbb{E}\left[Y+z(Y)\right]=\operatorname*{max}\left(\frac{1}{f_{\operatorname*{max}}},\frac{\mathbb{E}[(Y+z(Y))^{2}]}{2\beta}\right)$ (62)
    with
    $z(Y)=\max(0,\beta-Y).$ (63)
    [63, Algorithm 2] presents a low-complexity algorithm to solve $\beta$.
VIII-A2 Remote Decision Making
We consider the following decision making policies:
  • Co-design with goal-oriented sampling: This decision-making policy is obtained by solving Problem 1. For conciseness, we refer to the co-design approach simply as “goal-oriented sampling”.
  • Myopic decision policy $\pi_{\text{myopic}}$: At each time step, the remote decision maker selects an action based on the most recent (possibly outdated) observation $X_{t-\Delta(t)}$. The selected action aims to minimize the instantaneous cost and is formally given by
    $a_{t}=\pi_{\text{myopic}}(X_{t-\Delta(t)})=\arg\min_{a}\mathcal{C}(X_{t-\Delta(t)},a),$ (64)
    where $\mathcal{C}(\cdot,\cdot)$ denotes the one-step cost function. This policy is termed myopic because it excludes any prediction or planning mechanisms; instead, decisions are made greedily to minimize the one-step cost based on the most recently available observation.
  • Long-term optimal decision policy $\pi^{\star}$: The long-term optimal policy thinks ahead to minimize the long-term average cost over time. We compute it by solving the ACOE of the primal MDP using the RVI algorithm. The resulting policy employed at the remote decision maker $\pi^{\star}:\mathcal{S}\rightarrow\mathcal{A}$ selects an action based on the latest available observation $X_{t-\Delta(t)}$, according to $a_{t}=\pi^{\star}(X_{t-\Delta(t)})$. In our experiments and analysis, we use this decision-making policy as the default for the benchmark sampling policies.
Figure 7: Average cost vs. Average Delay $\mathbb{E}[Y_{i}]$.
Figure 8: Average cost vs. Maximum sampling frequency $f_{\max}$.
VIII-B Simulation Parameter Setup
VIII-B1 Primal MDP
$s_{0}$$s_{1}$$a_{0}:0.1$$a_{0}:0.1$$a_{0}:0.9$$a_{0}:0.9$$a_{1}:0.4$$a_{1}:0.01$$a_{1}:0.6$$a_{1}:0.99$
Figure 9: Simulation Setup: Transition diagram of the primal MDP given in Appendix LABEL:appendixh.
As a case study, we specifically consider a benchmark setup for clarity and insight. In this setup, the sizes of the state space and the action space of the primal MDP are both $2$, consisting of states $s_{0},s_{1}$ and actions $a_{0},a_{1}$ [^12] The binary state space is broadly representative of many real-world systems. For instance, it can model the occurrence or absence of critical events such as fires, industrial failures, security breaches, or anomalies in equipment status. The binary action space can be interpreted as whether or not to apply a control intervention in response to the critical event, e.g. extinguishing a detected fire or repairing a malfunctioning piece of equipment.. The primal MDP tuple is detailed in Appendix LABEL:appendixh, and is visualized by transition diagram given in Fig. 9.
In the primal MDP, the corresponding myopic decision-making policy is
$\pi_{\text{myopic}}(s_{0})=a_{0},\pi_{\text{myopic}}(s_{1})=a_{0},$ (65)
indicating a constant action regardless of state. Consequently, this decision making policy does not utilize state information at the remote decision maker, and thus the value of information transmission is null in this context.
The long-term optimal decision policy is solved by using RVI algorithm, and the solution is given as:
$\pi^{\star}(s_{0})=a_{1},\pi^{\star}(s_{1})=a_{0},$ (66)
which clearly depends on the system state. In this case, the remote decision maker must rely on the potentially stale state to select the “right” action, highlighting the utility of information transmission under the decision making policy.
VIII-B2 Random Delay Distribution
As a representative example, we consider an i.i.d. Bernoulli random delay model in which each delay $Y_{i}$ takes value $1$ with probability $p$ and $11$ with probability $1-p$, i.e., $\Pr(Y_{i}=1)=p,\Pr(Y_{i}=11)=1-p,\forall i\in\mathbb{Z}^{+}$. This binary-delay formulation enables tractable analysis while capturing key aspects of stochastic delay. The framework, however, is not limited to this case and can be generalized to arbitrary delay distributions.
VIII-C Discussions and Insights
Fig. 7 illustrates the performance comparison between various sampling and decision-making policies under different average delays $\mathbb{E}[Y_{i}]$. The proposed goal-oriented sampling consistently outperforms all benchmark methods across the full range of average delay values. A key observation is that the myopic decision-making policy, which always selects a fixed action regardless of state information, yields a constant cost independent of delay. This highlights an insight: in scenarios where the control policy is insensitive to state information, enhancing the communication channel (e.g., by reducing delay) does not lead to any improved goal-oriented decision making performance.
Moreover, Fig. 7 demonstrates that the performance gap between the proposed goal-oriented sampling policy and the myopic decision-making policy gradually narrows as the average delay $\mathbb{E}[Y_{i}]$ increases. This convergence reflects a fundamental limitation in decision-making under communication constraints: when the delay becomes excessively large, the received state information is so outdated that it loses its relevance to the current system condition. As a result, the optimal goal-oriented decision making degenerates into state-independent behavior when the average delay is large. In such regimes, the value of state information diminishes, and the benefits of optimized information transmission vanish.
Furthermore, Fig. 7 demonstrates that when the average delay is relatively small, AoI-optimal, zero-wait, and constant-wait sampling strategies also perform competitively, approaching the performance of goal-oriented sampling. However, their effectiveness deteriorates rapidly under large delays, sometimes falling below that of the constant-action myopic policy. This occurs because when delays are minimal, the received state information is still timely enough to inform near-optimal decisions. Conversely, under large average delay, the received data becomes too stale to offer meaningful advantage for decision making.
Fig. 8 illustrates the relationship between average cost and the maximum sampling frequency $f_{\max}$. The proposed goal-oriented sampling consistently achieves the lowest cost across all values of $f_{\max}$. As $f_{\max}$ increases, the performance of both goal-oriented sampling and AoI-optimal sampling improves gradually. Notably, the goal-oriented sampling curve aligns with the sensitivity analysis presented in Theorem 9, which predicts that the average cost initially decreases with $f_{\max}$, then saturates to a constant value.
In contrast, the uniform sampling policy in Fig. 8 exhibits a U-shaped cost curve. As $f_{\max}$ decreases, its performance converges to that of AoI-optimal sampling policy. This convergence highlights a key insight: under sparse sampling constraints, the transmitted information becomes too outdated/uninformative to support effective state-dependent decision-making, resulting in uniformly poor performance. Conversely, when the sampling interval $d=1/f_{\max}$, the system begins to accumulate a queuing backlog, resulting in increased delivery delays. This queuing-induced staleness reduces the timeliness of information at the decision-maker, thereby degrading goal-oriented decision-making performance.
IX Conclusion
In this paper, we have proposed a new remote MDP problem, namely AR-MDP in the time-lag MDP framework. Specifically, AoI, typically an optimization indicator for information freshness, is incorporated into AR-MDP both as a controllable random processing delay and as critical side information to support remote decision making. To investigate the fundamental trade-off between communication constraint and the decision-making effectiveness, we considered both sampling frequency constraint and random delay in this problem and developed low-complexity one-layer algorithms, OnePDSI and QuickBLP to solve this problem efficiently. Through theoretical analysis and experiments, we reveal how communication-induced information staleness and the resulting uncertainty can degrade the value of information and negatively impact remote decision-making performance.
Acknowledgment
By extending [1], this study further addresses the practical constraint of limited sampling frequency and proposes two low-complexity one-layer algorithms: OnePDSI and QuickBLP to address the sampling-constrained AR-MDP problem. The convergence of both methods is rigorously established in this paper. Beyond these developments, the analysis offers deeper insights into the interplay between communication delay, information freshness, and goal-oriented decision-making. We are grateful to Prof. Elif Uysal for her insightful discussions, and we also thank the reviewers of IEEE ITW 2024 for their valuable feedback, which has helped improve this work.